OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the second equation is substituted into the first equation. 2x – 7y = 4 3x + y = –17

OpenStudy (anonymous):

So I'd solve for one variable in one equation. Let's go with the second one. That'd be y=-17-3x. Now, plug that in for y in the original equation: 2x - 7(-17-3x) = 4 2x+21x+119=4 23x+119=4 23x=-115 x=-5 Now you have x. Substitute -5 for x in to any equation. Ill choose the second one. 3(-5)+y=-17 -15+y=-17 y=-2 And there you have it. The answer to this systems of equations is (-5, -2). Hope that helps!

OpenStudy (anonymous):

so out of these which one would it be? 2(3x – 17) – 7y = 4 2(–3x – 17) – 7y = 4 2x – 7(3x – 17) = 4 2x – 7(–3x – 17) = 4

OpenStudy (anonymous):

Well, here's a hint. The choices show the equation when y is substituted. What do you think?

OpenStudy (anonymous):

the last one?

OpenStudy (anonymous):

Yes, that's correct!